A 0.187 M weak acid solution has a pH of 3.99. Find Ka for the acid.
Express your answer using two significant figures.
1
Expert's answer
2020-07-10T05:25:33-0400

A 0.187 M weak acid solution has a pH of 3.99.


pH of 3.99. = 4 ( approx ) .


[H+] = 10-4 M .


A 0.187 M weak acid solution = C .


CONCENTRATION OF H+ = UNDER ROOT ( C* Ka ) . .


So , Ka = [H+]2 / C . .


So , Ka = 10-8 / 0.187 . = ( 5.347 * 10 -8 ) .( Approx . ) . Answer .........